# IncEventGS: Pose-Free Gaussian Splatting from a Single Event Camera

### v1

@@ 3.3 Camera Motion Trajectory Modeling
Since each event chunk usually contains too many events, we sample a portion of them according to the total number of events during optimization. We formulate the corresponding poses (i.e. $\mathbf{T}_{k}$ and $\mathbf{T}_{k+\Delta t}$) at the beginning and end of the sampled event portion within each chunk, by employing a camera motion trajectory. The trajectory is represented through linear interpolation between two camera poses, one at the beginning of the chunk $\mathbf{T}_{\mathrm{start}}\in\mathbf{SE}(3)$ and the other at the end $\mathbf{T}_{\mathrm{end}}\in\mathbf{SE}(3)$. The camera pose at time $t_{k}$ can thus be expressed as follows:
@@
where $t_{start}$ and $t_{end}$ represent the timestamps corresponding to the boundary of the event chunk. It follows that $\mathbf{T}_{k}$ is differentiable with respect to both $\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$. The objective of IncEventGS is thus to estimate both $\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$ for each event chunk, along with the learnable parameters of 3D Gaussians $\mathbf{G}_{\theta}$.
@@ 3.4 Incremental Tracking and Mapping
For both tracking and mapping, we exploit the previously introduced event data formation model to compute the loss from the synthesized and real accumulated event images. In particular, we compute the loss of the latest event chunk only for the tracking stage and minimize the following energy function:
@@
where both $\mathbf{E}(\mathbf{x})$ and $\hat{\mathbf{E}}(\mathbf{x})$ are the accumulated real and synthesized event images respectively, corresponding to a randomly sampled event portion within the latest event chunk.
Once the tracking is done, we insert the latest event chunk to mapper and exploit the estimated $\mathbf{T}_{start}^{*}$ and $\mathbf{T}_{end}^{*}$ as the initial value of the chunk to perform dense bundle adjustment. For computational consideration, we exploit a sliding window BA of the latest $n_{w}$ chunks and $n_{w}$ is a hyperparameter. In particular, we optimize both the motion trajectories and the 3D-GS jointly by minimizing the following loss functions:
@@
$\mathcal{L}_{event}$ $=\frac{1}{n_{w}}\sum_{i=0}^{n_{w}}\left\lVert\mathbf{E}_{i}(\mathbf{x})-\hat{\mathbf{E}}_{i}(\mathbf{x})\right\rVert_{2},$ (12)
$\mathcal{L}_{ssim}$ $=\frac{1}{n_{w}}\sum_{i=0}^{n_{w}}SSIM(\mathbf{E}_{i}(\mathbf{x}),\hat{\mathbf{E}}_{i}(\mathbf{x}))$ (13)
where $\lambda$ is a hyperparameter, SSIM is the structural dissimilarity loss Wang et al. (2004), both $\mathbf{E}_{i}(\mathbf{x})$ and $\hat{\mathbf{E}}_{i}(\mathbf{x})$ are the corresponding accumulated real and synthesized event images of the latest $i^{th}$ event portion respectively. As the event data streams in, we alternatively perform tracking and mapping.
3D-GS Initialization and System Boot-strapping. We initialize the 3D-GS by sampling point cloud randomly within a bounding box. The first $m$ event chunks (where $m$ is a hyperparameter) are selected for initialization, and all corresponding camera poses (e.g., $\mathbf{T}_{*}$) are randomly initialized to be near the identity matrix. We then minimize the loss computed by Eq. (11) with respect to the attributes of 3D-GS and the parameters of camera motion trajectories jointly.
Through experiments, we found that the above initialization procedure consistently produces high-quality brightness images. However, the 3D structure remains of low quality due to the short baselines of the event chunks. We further find that it could potentially affect the performance of the whole pipeline as more event data is received. Therefore, we utilize a monocular depth estimation network Ke et al. (2024) to predict a dense depth map from the rendered brightness image. This depth map is then used to re-initialize the centers of the 3D Gaussians by unprojecting the pixel depths, after which we repeat the minimization of Eq. (11) for system bootstrapping.
3D-GS Incrementally Growing. As the camera moves, new Gaussians is periodically introduced to cover newly explored regions. After tracking, we obtain an accurate camera pose estimate for each new event chunk. The center of new Gaussians are determined by:
$p=T\cdot\pi^{-1}(u,d_{u})$ (14)
where $p\in\mathbb{R}^{3}$ is a 3D point, $u\in\mathbb{R}^{2}$ is a point in the image plane, $d_{u}$ is depth of the 3D point $p$ projecting on pixel u, which is rendered by equation 5, $\pi^{-1}$ denotes camera inverse projection, $T$ is the camera pose from tracking. To ensure that new Gaussians are only added in previously unmapped areas, a visibility mask is generated to guide the expansion of the Gaussian splatting process, as following:
$M(p)=V<\lambda_{V}$ (15)
where $V$ is the rendered alpha map, $\lambda_{V}$ is the hyperparameter.
@@ 4 Experiments
All experiments were conducted on a desktop PC equipped with a 5.73GHz AMD Ryzen 9 7900x CPU and an NVIDIA RTX 3090 GPU. The first $m=3$ event chunks were used for initialization. During the mapping stage, a sliding window size of $n_{w}=20$ was employed for the bundle adjustment algorithm. The hyperparameters were set as follows: $\lambda=0.05$, $\lambda_{V}=0.8$, and $n_{seg}=100$. For the synthetic dataset, $n_{low}=400k$ and $n_{up}=500k$, while for the real dataset, $n_{low}=60k$ and $n_{up}=80k$. Each event chunk had a time interval of 50 ms. The learning rate of the camera poses is set to 1e-4 and that for the attributes of 3D-GS are set the same as the original 3D-GS work. The number of optimization steps for initialization is 4500, and that for tracking and mapping are set to 200 and 1500 respectively. The contrast threshold $C$ of the event camera is set to 0.1 for synthetic datasets and 0.2 for real datasets empirically.
Baselines and Evaluation Metrics. IncEventGS performs incremental dense 3D-GS reconstruction and motion estimation using only a single event camera. To the best of our knowledge, there are no existing event-only NeRF or 3D-GS SLAM methods that do not rely on ground-truth poses, making direct comparisons challenging. Therefore, we conduct a thorough comparison of our method with several event-based NeRF approaches, including E-NeRF Klenk et al. (2022), EventNeRF Rudnev et al. (2023a), and Robust e-NeRF Low & Lee (2023), as well as two-stage methods such as E2VID Rebecq et al. (2019) + COLMAP Schönberger & Frahm (2016) + 3DGS Kerbl et al. (2023), and E2VID Rebecq et al. (2019) + DEVO Klenk et al. (2024) + 3DGS Kerbl et al. (2023). E-NeRF, EventNeRF, and Robust e-NeRF leverage implicit neural radiance fields for 3D scene representation, requiring ground-truth camera poses for accurate NeRF reconstruction. The two-stage methods we examine include E2VID + COLMAP + 3DGS and E2VID + DEVO + 3DGS. In those approaches, event data is first converted into grayscale images using E2VID. The subsequent steps differ: in the E2VID + COLMAP + 3DGS method, camera poses are estimated from these images using COLMAP, while in the E2VID + DEVO + 3DGS method, poses are estimated using DEVO. Finally, 3D-GS is trained with the generated images and poses in both methods. Both the quantitative and qualitative comparisons are performed on the synthetic dataset. Since there are no paired ground truth images for the real dataset, we only perform qualitative comparisons on the real dataset. In terms of motion trajectory evaluations, we use the publicly available state-of-the-art event-only visual odometry method, i.e. DEVO Klenk et al. (2024), for comparison, both quantitatively and qualitatively.
The metrics used for novel view synthesis (NVS) include the commonly employed PSNR, SSIM, and LPIPS. For motion trajectory evaluations, we utilize Absolute Trajectory Error (ATE). To ensure fair comparisons, we employ the evaluation code provided by EventNeRF to compute the NVS metrics, which applies a linear color transformation between predictions and ground truth. Additionally, we use the public EVO toolbox Grupp (2017) to compute the trajectory metrics.
@@
To properly evaluate the performance of NVS and motion trajectory estimation, we synthesized event data using the 3D scene models from the Replica dataset Straub et al. (2019). In particular, we exploit the room0, room2, office0, office2, and office3 scenes. We rendered high frame rate RGB images at 1000 Hz with a resolution of 768x480 pixels. Those images are then converted to grayscale and the event data is generated via the events simulator Gehrig et al. (2020). The contrast threshold is set to 0.1. To simulate real-world camera motions, we exploit the same motion trajectories as that of NICE-SLAM Zhu et al. (2022) for data generation.
We used the event dataset provided by TUM-VIE Klenk et al. (2021) for real data evaluations, which is also used by E-NeRF and Robust e-NeRF. TUM-VIE Klenk et al. (2021) captured the event datasets by a pair of Prophesee Gen4 HD event cameras with a resolution of 1280x720 pixels. We only use the left-event camera data for our experiment.
4.2 Quantitative Evaluations.
We conduct quantitative evaluations against event NeRF methods(E-NeRF, EventNeRF, and Robust e-NeRF) and two-stage methods (E2VID + COLMAP + 3DGS and E2VID + DEVO + 3DGS) in terms of the quality of NVS and pose estimation performance.
The NVS performance is evaluated on Replica-dataset and the results are presented in Table 1. It is important to note that the metrics are lower than those typically observed in standard NeRF/3D-GS methods for RGB images, primarily due to the absence of adequate RGB image supervision. Even though nerf-based methods use ground truth poses for training, IncEventGS still significantly outperforms them, highlighting the advantages of our approach utilizing a 3D Gaussian representation. Additionally, our method greatly surpasses two-stage methods that also employ 3D Gaussian representation, demonstrating superior pose estimation and the effectiveness of our bundle adjustment technique.
Table 1: NVS performance comparison on Replica dataset. The result demonstrates that our method outperforms NeRF-based and two-stage methods.
@@
Table 2: Pose accuracy (ATE, cm) on Replica and TUM-VIE datasets. The results demonstrate that our method delivers better performance in terms of camera motion estimation.
@@
We evaluate pose estimation performance using the $ATE$ metric on both synthetic and real datasets, comparing our method with EVO and E2VID + COLMAP. The results, presented in Table 2, show that our method outperforms both baselines, validating the effectiveness of our incremental tracking and mapping technique.
@@
4.3 Qualitative Evaluations.
We evaluate our method against event NeRF methods and two-stage methods qualitatively in terms of novel view image synthesis, both on synthetic and real data. The results are presented in both Fig. 2 and Fig. 3. It demonstrates that our method can deliver better novel view images, while event NeRF methods and two-stage methods render images with additional artifacts. Compared to NeRF-based methods, our approach demonstrates the advantage of IncEventGS by leveraging 3D Gaussian Splatting as the underlying scene representation. In contrast to two-stage methods, our dense bundle adjustment optimizes both 3D Gaussian Splatting and camera pose using event data, whereas two-stage approaches tend to accumulate errors over time, as confirmed by the experimental results. We also provide representative visualization of ATE error mapped onto trajectories in Fig. 4, both on synthetic and real dataset. It demonstrates that IncEventGS is able to recover more accurate motion trajectories.
4.4 Ablation Study
We conduct ablation studies to confirm our design choices. In particular, we study the effect of a monocular depth estimation network for system bootstrapping and event slicing hyperparameters $n_{low}$, $n_{up}$. The experiments are conducted with the Replica dataset and the results are shown in Table 4 and Table 4 respectively.
We found that depth initialization significantly impacts pose estimation, reducing the Average Trajectory Error (ATE) from 1.534 cm to 0.064 cm. Additionally, this improvement in pose estimation leads to a slight enhancement in Novel View Synthesis (NVS) performance. These results verify the importance of using depth initialization.
We compare several combinations of hyperparameters $n_{\text{low}}$ and $n_{\text{up}}$, which refer to the range of event slicing window sizes. Table 4 demonstrates that both too small and too large window sizes negatively impact the performance of Novel View Synthesis (NVS) and pose estimation. Consequently, we select $n_{\text{low}}=400k$ and $n_{\text{up}}=500k$ for our experiments on the Replica dataset.
5 Conclusion
We present the first incremental 3D dense reconstruction algorithm, i.e. IncEventGS, with a single event camera under the framework of 3D Gaussian Splatting. We adopt the tracking and mapping paradigm in conventional SLAM pipeline to do incremental motion estimation and 3D scene reconstruction simultaneously. To handle the continuous and asynchronous characteristics of event stream, we exploit a continuous trajectory model to model the event data formation process. The experimental results on both synthetic and real datasets demonstrate the superior performance of IncEventGS over prior state-of-the-art methods in terms of high-quality novel image synthesis and camera pose estimation.
room0 room2 office0 office2 office3
E2VID+ DEVO+3DGS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 5 Conclusion
E2VID+ DEVO+3DGS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 3: Qualitative evaluation for novel view image synthesis on real dataset. It demonstrates that our method is able to render better images with fewer artifacts than event NeRF methods and two-stage methods. Note there are no GT images aligned with the event camera, we choose closest images of RGB camera and crop it to the the size with rendered images for visual comparisons.

### v4

@@ 3.3 Camera Motion Trajectory Modeling
Since each event chunk usually contains too many events, we sample a portion of them according to the total number of events during optimization. Following [38, 32], we formulate the corresponding poses (i.e. $\mathbf{T}_{k}$ and $\mathbf{T}_{k+\Delta t}$) at the beginning and end of the sampled event portion within each chunk, by employing a camera motion trajectory. The trajectory is represented through linear interpolation between two camera poses, one at the beginning of the chunk $\mathbf{T}_{\mathrm{start}}\in\mathbf{SE}(3)$ and the other at the end $\mathbf{T}_{\mathrm{end}}\in\mathbf{SE}(3)$. The camera pose at time $t_{k}$ can thus be expressed as follows:
@@
where $t_{start}$ and $t_{end}$ represent the timestamps corresponding to the boundary of the event chunk. It follows that $\mathbf{T}_{k}$ is differentiable with respect to both $\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$. The objective of IncEventGS is thus to estimate both $\mathbf{T}_{\mathrm{start}}$ and $\mathbf{T}_{\mathrm{end}}$ for each event chunk, along with the learnable parameters of 3D-GS $\mathbf{G}_{\theta}$.
@@ 3.4 Incremental Tracking and Mapping
For both tracking and mapping, we aim to minimize the difference between the synthesized and measured brightness changes. In particular, we compute the loss of the latest event chunk only for the tracking stage and minimize the following loss function:
@@
where $\hat{\mathbf{E}}(\mathbf{x})$ and $\mathbf{E}(\mathbf{x})$ are the synthesized and measured brightness changes respectively, corresponding to a randomly sampled event portion within the latest event chunk.
Once the tracking is done, we insert the latest event chunk to the mapper and exploit the estimated $\mathbf{T}_{start}^{*}$ and $\mathbf{T}_{end}^{*}$ as the initial value of the chunk to perform dense bundle adjustment. For computational consideration, we exploit a sliding window BA of the latest $n_{w}$ chunks, and $n_{w}$ is a hyperparameter. In particular, we optimize both the motion trajectories and the 3D-GS jointly by minimizing the following loss functions:
@@
$\mathcal{L}_{event}$ $=\left\lVert\mathbf{E}_{i}(\mathbf{x})-\hat{\mathbf{E}}_{i}(\mathbf{x})\right\rVert_{2},$ (12)
$\mathcal{L}_{ssim}$ $=SSIM(\mathbf{E}_{i}(\mathbf{x}),\hat{\mathbf{E}}_{i}(\mathbf{x})),$ (13)
where $\lambda$ is a hyperparameter, SSIM is the structural dissimilarity loss [34]. As the event data streams in, we alternatively perform tracking and mapping.
3.5 System Initialization and Boot-strapping
Conventional frame-based 3D-GS methods usually require a good initial point cloud and camera poses computed via COLMAP [27] for initialization. However, they are usually not easy to obtain if we are only given a single event camera. We therefore initialize the 3D-GS by sampling point cloud randomly within a bounding box. The first $m$ event chunks (where $m$ is a hyperparameter) are selected for initialization, and all corresponding camera poses are randomly initialized to be near the identity matrix. We then minimize the loss computed by Eq. (11) with respect to the attributes of 3D-GS and the parameters of camera motion trajectories jointly.
Through experiments, we found that the above initialization procedure consistently produces satisfactory brightness images. However, the 3D structure remains of low quality due to the short baselines of the moving event camera. We further find that it could potentially affect the performance of the whole pipeline without a good initial 3D structure as more event data is received. Therefore, we utilize a monocular depth estimation network [9] to predict a dense depth map from the rendered brightness image after the pipeline is trained for certain iterations. This depth map is then used to re-initialize the centers of the 3D Gaussians by unprojecting the pixel depths, after which we repeat the minimization of Eq. (11) for system bootstrapping. More details about system re-initialization can be found in our supplementary material.
3D-GS Incrementally Growing. As the camera moves, new Gaussians is periodically introduced to cover newly explored regions. After tracking, we obtain an accurate camera pose estimate for each new event chunk. The centers of new Gaussians are determined by:
$\mathbf{p}=\mathbf{T}\cdot\pi^{-1}(\mathbf{u},d_{u})$ (14)
where $\mathbf{u}\in\mathbb{R}^{2}$ is pixel coordinate in the image plane, $d_{u}$ is depth of the 3D point $\mathbf{p}$ projecting onto the image plane, which is rendered by Eq. (5), $\pi^{-1}$ denotes camera inverse projection, $\mathbf{T}$ is the camera pose from tracking. To ensure that new Gaussians are only added in previously unmapped areas, a visibility mask is computed to guide the expansion of the Gaussian splatting process, as following:
$\mathbf{M}(p)=V<\lambda_{V},$ (15)
where $\mathbf{V}$ is the rendered alpha map and $\lambda_{V}$ is a hyperparameter.
@@ 4 Experiments
All experiments are conducted on a desktop PC equipped with a 5.73GHz AMD Ryzen 9 7900x CPU and an NVIDIA RTX 3090 GPU. The first $m=3$ event chunks are used for initialization. During the mapping stage, a sliding window size of $n_{w}=20$ is employed for the bundle adjustment algorithm. The hyperparameters are set as follows: $\lambda=0.05$, $\lambda_{V}=0.8$, and $n_{seg}=100$. For the synthetic dataset, $n_{low}=400k$ and $n_{up}=500k$, while for the real dataset, $n_{low}=60k$ and $n_{up}=80k$. Each event chunk has a time interval of 50 ms. The learning rate of the camera poses is set to 1e-4 and that for the attributes of 3D-GS are set the same as the original 3D-GS work. The number of optimization steps for initialization is 4500, and that for tracking and mapping are set to 200 and 1500 respectively. The contrast threshold $C$ of the event camera is set to 0.1 for synthetic datasets and 0.2 for real datasets empirically.
Baselines and Evaluation Metrics. To the best of our knowledge, there are no existing event-only NeRF or 3D-GS methods that do not rely on ground-truth poses, making direct comparisons challenging. Therefore, we conduct a thorough comparison of our method with several event-based NeRF approaches, including E-NeRF [14], EventNeRF [25], and Robust e-NeRF [17], as well as our custom implemented two-stage method (i.e. E2VID [24] + COLMAP [27] + 3DGS [11]). E-NeRF, EventNeRF, and Robust e-NeRF leverage implicit neural radiance fields for 3D scene representation, requiring ground-truth camera poses for accurate NeRF reconstruction. For E2VID + COLMAP + 3DGS, event data is first converted into brightness images using E2VID. The camera poses are then estimated from these images using COLMAP, and 3D-GS is trained with the generated images and poses. Both the quantitative and qualitative comparisons are performed on the synthetic dataset. Since there are no paired ground truth images for the real dataset, we only perform qualitative comparisons on the real dataset. In terms of motion trajectory evaluations, we use the publicly available state-of-the-art event-only visual odometry methods, i.e. DEVO (mono) [15] and ESVO2(stereo) [21], for comparison.
room0 room2 office0 office2 office3
The metrics used for novel view synthesis (NVS) include the commonly employed PSNR, SSIM, and LPIPS. For motion trajectory evaluations, we utilize Absolute Trajectory Error (ATE). To ensure fair comparisons, we employ the evaluation code provided by EventNeRF to compute the NVS metrics, which applies a linear color transformation between predictions and ground truth. Additionally, we use the public EVO toolbox [5] to compute the trajectory metrics.
@@
To properly evaluate the performance of NVS and motion trajectory estimation, we synthesized event data using the 3D scene models from the Replica dataset [28]. In particular, we exploit the room0, room2, office0, office2, and office3 scenes. We render high frame rate RGB images at 1000 Hz with a resolution of 768x480 pixels. These images are then converted to grayscale, and the event data is generated via the events simulator [4]. The contrast threshold is set to 0.1. To simulate real-world camera motions, we exploit the same motion trajectories as that of NICE-SLAM [39] for data generation.
We use the event dataset provided by TUM-VIE [13] for real data evaluations, which is also used by E-NeRF and Robust e-NeRF. TUM-VIE captures the event datasets by a pair of Prophesee Gen4 HD event cameras with a resolution of 1280x720 pixels. We only use the left-event camera data for our experiment.
@@
Table 1: NVS performance comparison on Replica dataset. All results are calculated using the same code and ground-truth images. The results demonstrate that our method outperforms NeRF-based and two-stage methods.
@@
Table 2: Pose accuracy (ATE, cm) on Replica and TUM-VIE datasets. The results demonstrate that our method delivers better performance in terms of camera motion estimation. Since ESVO2 is a stereo event camera method, it cannot run on the Replica dataset.
4.2 Ablation Study
@@
We conduct ablation studies to confirm our design choices. In particular, we study the effect of a monocular depth estimation network for system bootstrapping and event slicing hyperparameters $n_{low}$, $n_{up}$. The experiments are conducted with the Replica dataset, and the results are shown in Table 3 and Table 4, respectively.
We found that depth initialization significantly impacts pose estimation, reducing the Average Trajectory Error (ATE) from 1.534 cm to 0.064 cm. Additionally, this improvement in pose estimation leads to a slight enhancement in Novel View Synthesis (NVS) performance. These results verify the importance of using depth initialization during the boot-strapping stage.
We compare several combinations of hyperparameters $n_{\text{low}}$ and $n_{\text{up}}$, which refer to the range of event slicing window size. Table 4 demonstrates that both too small and too large window sizes negatively impact the performance of Novel View Synthesis (NVS) and pose estimation. Consequently, we select $n_{\text{low}}=400k$ and $n_{\text{up}}=500k$ for our experiments on the Replica dataset.
4.3 Quantitative Evaluations.
We conduct quantitative evaluations against event NeRF methods(E-NeRF, EventNeRF, and Robust e-NeRF) and our custom implemented two-stage method (i.e. E2VID + COLMAP + 3DGS) in terms of the quality of NVS and pose estimation performance.
The NVS performance is evaluated on Replica-dataset and the results are presented in Table 1. It is important to note that the metrics are lower than those typically observed in standard NeRF/3D-GS methods for RGB images, primarily due to the lack of absolute brightness supervision. Even though NeRF-based methods use ground truth poses for training, IncEventGS still significantly outperforms them, highlighting the advantages of our approach utilizing a 3D Gaussian representation. Additionally, our method greatly surpasses two-stage method that also employs 3D Gaussian representation, demonstrating superior pose estimation and the effectiveness of our bundle adjustment technique.
We evaluate pose estimation performance using the ATE metric on both synthetic and real datasets, comparing our method with DEVO, ESVO2 and E2VID + COLMAP. The results, presented in Table 2, show that our method outperforms both baselines, validating the effectiveness of our incremental tracking and mapping technique.
@@
Figure 3: Qualitative evaluation for novel view image synthesis on real dataset. It demonstrates that our method is able to render better images with fewer artifacts than event NeRF methods and two-stage methods. Note that there are no GT images aligned with the event camera, and we choose the closest images from the RGB camera and crop them to the same size as the rendered images for visual comparisons.
@@
4.4 Qualitative Evaluations.
We evaluate our method against event NeRF methods and two-stage method qualitatively in terms of novel view image synthesis, both on synthetic and real data. The results are presented in both Fig. 2 and Fig. 3. It demonstrates that our method can deliver better novel view images, while event NeRF methods and two-stage method render images with additional artifacts. Compared to NeRF-based methods, our approach demonstrates the advantage of IncEventGS by leveraging 3D Gaussian Splatting as the underlying scene representation. In contrast to two-stage method, our dense bundle adjustment optimizes both 3D Gaussian Splatting and camera pose using event data, whereas two-stage approaches tend to accumulate errors over time, as confirmed by the experimental results. We also provide representative visualization of ATE error mapped onto trajectories in Fig. 4, both on synthetic and real dataset. It demonstrates that IncEventGS is able to recover more accurate motion trajectories.
5 Conclusion
We present the first pose-free 3D Gaussian splatting reconstruction model based on event camera, i.e. IncEventGS, and is useful for real world applications. We adopt the tracking and mapping paradigm in conventional SLAM pipeline to do incremental motion estimation and 3D scene reconstruction simultaneously. To handle the continuous and asynchronous characteristics of event stream, we exploit a continuous trajectory model to model the event data formation process. The experimental results on both synthetic and real datasets demonstrate the superior performance of IncEventGS over prior state-of-the-art methods in terms of high-quality novel image synthesis and camera pose estimation.
Acknowledgements. This work was supported in part by NSFC under Grant 62202389, in part by a grant from the Westlake University-Muyuan Joint Research Institute, and in part by the Westlake Education Foundation.
Supplementary Material
Figure 5: The re-initialization process of IncEventGS.
6 More Details about Re-initialization
The re-initialization process is illustrated in Fig. 5. After the first-time initialization, we can render a brightness image from 3D-GS at pose $\mathbf{T}_{1}$, where $\mathbf{T}_{1}$ represents the camera pose at the end of the first event chunk. To improve the 3D structure of 3D-GS, we use a monocular depth estimation network [9] to predict a dense depth map from the rendered brightness image. This depth map is then used to re-initialize the centers of the 3D Gaussians by unprojecting the pixel depths at camera pose $\mathbf{T}_{1}$, as illustrated in Fig. 5. After re-centering the 3D Gaussians, we perform the initialization process again to achieve both accurate 3D structure and exceptional brightness image rendering performance.
7 Comparison with Gaussian-based Event Methods
To further evaluate our method, we conducted additional comparisons against state-of-the-art Gaussian-based event approaches. Since Event3DGS [35] had not been open-sourced, we chose to compare against E2GS[1] and EvGGS[31]. In particular, we removed the supervision of blurred image in E2GS and exploited the pretrained weight of EvGGS for comparisons. As shown in Table 6, our method still outperforms those two baselines event though they used ground truth poses. Since EvGGS is a generalizable method based on a feed-forward network, it has limited generalization capability on unseen dataset.
8 Experiments in Fast-Motion Scenarios
Fast camera movement can induce motion blur, making it challenging to reconstruct the scene and estimate camera poses using RGB-based algorithms. We compare our event-based method with two state-of-the-art pose-free Gaussian SLAM implementations: MonoGS [18] (RGB modality) and SplaTAM [10] (RGBD modality). By leveraging the high temporal resolution of event cameras, our method experiences minimal performance degradation, even under fast motion. Additionally, it is more effective at preserving high-frequency information in the scene. As shown in Fig. 6, our approach delivers superior novel view synthesis results, particularly during rapid camera movement.
room0 room2 office0 office2 office3
SplaTAM Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MonoGS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Blur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative evaluation of novel view image synthesis on the Replica dataset. The experimental results demonstrate that our method renders higher-quality images when the camera is moving fast.
9 Experiments on Color Event Datasets
Our method can also be applied to color event datasets by integrating the Bayer filter [25], as shown below:
$\mathcal{L}_{event}$ $=\left\lVert\mathbf{F}\odot\mathbf{E}_{i}(\mathbf{x})-\mathbf{F}\odot\hat{\mathbf{E}}_{i}(\mathbf{x})\right\rVert_{2}$ (16)
$\mathcal{L}_{ssim}$ $=SSIM(\mathbf{F}\odot\mathbf{E}_{i}(\mathbf{x}),\mathbf{F}\odot\hat{\mathbf{E}}_{i}(\mathbf{x}))$ (17)
Furthermore, our method can be extended to incorporate training with ground-truth poses.
We conducted experiments on the EventNeRF dataset [25], which focuses on object reconstruction. Due to the dataset’s limited features, pose estimation is challenging; neither COLMAP nor DEVO can estimate camera poses on this dataset. As shown in Fig. 7, our method can still successfully optimize both the 3D scene and camera poses even without ground-truth poses, though it produces minor artifacts. When trained with ground-truth poses, our method achieves improved novel view synthesis, with fewer artifacts and sharper textures.
lego materials drums chair ficus
Ours(wo) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours(w) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT RGB Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Qualitative evaluation of novel view image synthesis on color event dataset. Ours (wo) denotes our method trained without ground-truth camera poses, while Ours (w) denotes the method trained with ground-truth camera poses.
Method Synthetic ($768\times 480$) Real-world ($1280\times 720$)
Training Storage Training Storage
ENeRF 12 hour 253M 12 hour 253M
EventNeRF 21 hour 14M 24 hour 14M
Robust e-NeRF 11 hour 745M 13 hour 745M
Ours 0.5 hour 65M 2 hour 55M
Table 5: Average model efficiency comparison.
room0 room2 office0 office2 office3
[…]
E2GS* 21.75 0.77 0.25 23.11 0.82 0.20 20.09 0.75 0.18 18.62 0.78 0.20 20.13 0.84 0.16
EvGGS 15.16 0.37 0.62 15.85 0.34 0.61 18.51 0.37 0.59 10.95 0.27 0.69 13.13 0.29 0.66
Table 6: NVS performance comparison on Replica dataset. * denotes we removed the supervision of blurred images from the original E2GS.The result demonstrates that our method outperforms those two baseline methods.
10 Time Evaluations
As shown in Table 5, our method has a significant advantage in training time compared to NeRF-based methods. Additionally, our method achieves an NVS rendering speed of approximately 500 FPS, whereas NeRF-based methods reach only about 0.5 FPS.
[…]